OpenStudy (anonymous):

Traci is getting a new hairstyle. Her options are: Length: short, long, medium Texture: curly, straight Color: Black, Brown, Blonde, Red How many different hairstyles can Traci make from the choices shown?

OpenStudy (hoblos):

3C1 * 2C1 * 4C1 = 3*2*4 = 24